# Chiral Heisenberg Gross-Neveu-Yukawa criticality: honeycomb vs. SLAC fermions

### v1

@@ Appendix A Hopping amplitude of SLAC Fermions
For even $L$ system sizes we have the momenta ${k=2\pi m/L}$, where ${m=-M,\ldots,M-1\in\mathbb{N}}$ with ${M=L/2}$ and we can use again use Eq. (23) to proceed with ${N=L/2+1}$ in the same way as for Eq. (25) to first compute
$\varepsilon^{\prime}(x)$ $=\frac{4\mathrm{i}\pi}{L^{2}}\;\mathrm{Im}\left[\frac{\left(\frac{L}{2}+1\right)\,\mathrm{e}^{\mathrm{i}\frac{2\pi}{L}\left(\frac{L}{2}+1\right)x}\left(1-\mathrm{e}^{\mathrm{i}\frac{2\pi}{L}x}\right)-\mathrm{e}^{\mathrm{i}\frac{2\pi}{L}x}\left(1-\mathrm{e}^{\mathrm{i}\frac{2\pi}{L}\left(\frac{L}{2}+1\right)x}\right)}{\left(1-\mathrm{e}^{\mathrm{i}\frac{2\pi}{L}x}\right)^{2}}\right]$
$=\frac{4\mathrm{i}\pi}{L^{2}}\;\mathrm{Im}\left[\frac{\left(\frac{L}{2}+1\right)\,\mathrm{e}^{\mathrm{i}\pi x}\,\mathrm{e}^{\mathrm{i}\frac{2\pi}{L}x}\,\mathrm{e}^{\mathrm{i}\frac{\pi}{L}x}\left(\mathrm{e}^{-\mathrm{i}\frac{\pi}{L}x}-\mathrm{e}^{\mathrm{i}\frac{\pi}{L}x}\right)-\mathrm{e}^{\mathrm{i}\frac{2\pi}{L}x}\left(1-\mathrm{e}^{\mathrm{i}\pi x}\,\mathrm{e}^{\mathrm{i}\frac{2\pi}{L}x}\right)}{\mathrm{e}^{\mathrm{i}\frac{2\pi}{L}x}\left(\mathrm{e}^{-\mathrm{i}\frac{\pi}{L}x}-\mathrm{e}^{\mathrm{i}\frac{\pi}{L}x}\right)^{2}}\right]$
$=\frac{4\mathrm{i}\pi}{L^{2}}\;\mathrm{Im}\left[\frac{\left(\frac{L}{2}+1\right)\,\mathrm{e}^{\mathrm{i}\pi x}\,\mathrm{e}^{\mathrm{i}\frac{\pi}{L}x}(-2\mathrm{i})\,\sin\left(\frac{\pi}{L}x\right)-\left(1-\mathrm{e}^{\mathrm{i}\pi x}\,\mathrm{e}^{\mathrm{i}\frac{2\pi}{L}x}\right)}{-4\,\sin\left(\frac{\pi}{L}x\right)^{2}}\right]$
$=\frac{4\mathrm{i}\pi}{L^{2}}(-1)^{x}\frac{\left(\frac{L}{2}+1\right)\,(-2)\,\cos\left(\frac{\pi}{L}x\right)\,\sin\left(\frac{\pi}{L}x\right)+2\sin\left(\frac{\pi}{L}x\right)\cos\left(\frac{\pi}{L}x\right)}{-4\,\sin\left(\frac{\pi}{L}x\right)^{2}}$
$=\frac{\mathrm{i}\pi}{L}(-1)^{x}\frac{\cos\left(\frac{\pi}{L}x\right)}{\sin\left(\frac{\pi}{L}x\right)}\;,$ (26)
@@
and then subtract from $\varepsilon^{\prime}(x)$ the contribution ${m=L/2}$ to avoid double counting
@@
$=\frac{\mathrm{i}\pi}{L}(-1)^{x}\frac{\cos\left(\frac{\pi}{L}x\right)}{\sin\left(\frac{\pi}{L}x\right)}-\frac{\pi}{L}\,(-1)^{x}$
$=\frac{\mathrm{i}\pi}{L}(-1)^{x}\left(\frac{\cos\left(\frac{\pi}{L}x\right)}{\sin\left(\frac{\pi}{L}x\right)}+\mathrm{i}\right)\;.$ (27)
@@ Appendix B Additional correlations ratios & finite size scaling
The CRs for the momenta $(\mathbf{b}_{1},0)$, $(\mathbf{b}_{1},\mathbf{b}_{2})$ and $(2\mathbf{b}_{1},0)$ close to the Dirac nodal point for the ${N=8}$ SLAC formulation and the honeycomb lattice are presented without fits in Fig. 16. Lines are guides to the eye only. Fit results to the CRs without scaling corrections on both lattices are shown in Fig. 17.
Figure 16: The correlation ratios for the moment $(\mathbf{b}_{1},0)$, $(\mathbf{b}_{1},\mathbf{b}_{2})$ and $(2\mathbf{b}_{1},0)$ away from the Dirac nodal point for (a) the ${N=8}$ SLAC formulation and (b) the honeycomb lattice.

### v2

@@ Appendix A Hopping amplitude of SLAC Fermions
For even $L$ system sizes we have the momenta $k=2\pi m/L$}, where $m=-M,\ldots,M-1\in\mathbb{N}$} with M=L/2{M=L/2} and we can use again use Eq. (23) to proceed with N=L/2+1{N=L/2+1} in the same way as for Eq. (25) to first compute
$\varepsilon^{\prime}(x)$ $=-\frac{4\mathrm{i}\pi}{L^{2}}\;\mathrm{Im}\left[\frac{\left(\frac{L}{2}+1\right)\,\mathrm{e}^{\mathrm{i}\frac{2\pi}{L}\left(\frac{L}{2}+1\right)x}\left(1-\mathrm{e}^{\mathrm{i}\frac{2\pi}{L}x}\right)-\mathrm{e}^{\mathrm{i}\frac{2\pi}{L}x}\left(1-\mathrm{e}^{\mathrm{i}\frac{2\pi}{L}\left(\frac{L}{2}+1\right)x}\right)}{\left(1-\mathrm{e}^{\mathrm{i}\frac{2\pi}{L}x}\right)^{2}}\right]$
$=-\frac{4\mathrm{i}\pi}{L^{2}}\;\mathrm{Im}\left[\frac{\left(\frac{L}{2}+1\right)\,\mathrm{e}^{\mathrm{i}\pi x}\,\mathrm{e}^{\mathrm{i}\frac{2\pi}{L}x}\,\mathrm{e}^{\mathrm{i}\frac{\pi}{L}x}\left(\mathrm{e}^{-\mathrm{i}\frac{\pi}{L}x}-\mathrm{e}^{\mathrm{i}\frac{\pi}{L}x}\right)-\mathrm{e}^{\mathrm{i}\frac{2\pi}{L}x}\left(1-\mathrm{e}^{\mathrm{i}\pi x}\,\mathrm{e}^{\mathrm{i}\frac{2\pi}{L}x}\right)}{\mathrm{e}^{\mathrm{i}\frac{2\pi}{L}x}\left(\mathrm{e}^{-\mathrm{i}\frac{\pi}{L}x}-\mathrm{e}^{\mathrm{i}\frac{\pi}{L}x}\right)^{2}}\right]$
$=-\frac{4\mathrm{i}\pi}{L^{2}}\;\mathrm{Im}\left[\frac{\left(\frac{L}{2}+1\right)\,\mathrm{e}^{\mathrm{i}\pi x}\,\mathrm{e}^{\mathrm{i}\frac{\pi}{L}x}(-2\mathrm{i})\,\sin\left(\frac{\pi}{L}x\right)-\left(1-\mathrm{e}^{\mathrm{i}\pi x}\,\mathrm{e}^{\mathrm{i}\frac{2\pi}{L}x}\right)}{-4\,\sin\left(\frac{\pi}{L}x\right)^{2}}\right]$
$=-\frac{4\mathrm{i}\pi}{L^{2}}(-1)^{x}\frac{\left(\frac{L}{2}+1\right)\,(-2)\,\cos\left(\frac{\pi}{L}x\right)\,\sin\left(\frac{\pi}{L}x\right)+2\sin\left(\frac{\pi}{L}x\right)\cos\left(\frac{\pi}{L}x\right)}{-4\,\sin\left(\frac{\pi}{L}x\right)^{2}}$
$=-\frac{\mathrm{i}\pi}{L}(-1)^{x}\frac{\cos\left(\frac{\pi}{L}x\right)}{\sin\left(\frac{\pi}{L}x\right)}\;,$ (26)
@@
and then subtract from $\varepsilon^{\prime}(x)$ the contribution m=L/2{m=L/2} to avoid double counting
@@
$=-\frac{\mathrm{i}\pi}{L}(-1)^{x}\frac{\cos\left(\frac{\pi}{L}x\right)}{\sin\left(\frac{\pi}{L}x\right)}-\frac{\pi}{L}\,(-1)^{x}$
$=-\frac{\mathrm{i}\pi}{L}(-1)^{x}\left(\frac{\cos\left(\frac{\pi}{L}x\right)}{\sin\left(\frac{\pi}{L}x\right)}+\mathrm{i}\right)\;.$ (27)
@@ Appendix B Additional correlations ratios & finite size scaling
The CRs for the momenta $(\mathbf{b}_{1},0)$, $(\mathbf{b}_{1},\mathbf{b}_{2})$ and $(2\mathbf{b}_{1},0)$ close to the Dirac nodal point for the N=8{N=8} SLAC formulation and the honeycomb lattice are presented without fits in Fig. 16. Lines are guides to the eye only. Fit results to the CRs without scaling corrections on both lattices are shown in Fig. 17.
Figure 16: The correlation ratios for the moment $(\mathbf{b}_{1},0)$, $(\mathbf{b}_{1},\mathbf{b}_{2})$ and $(2\mathbf{b}_{1},0)$ away from the Dirac nodal point for (a) the N=8{N=8} SLAC formulation and (b) the honeycomb lattice.